OpenStudy (anonymous):

A cone has volume where r is the radius of the base and h is the height. If a cone has a volume cm3 and the radius is 8 centimeters, find its height

OpenStudy (anonymous):

A cone has volume where r is the radius of the base and h is the height. If a cone has a volume 192picm3 and the radius is 8 centimeters, find its height

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

volume of a cone = 1/3 ( pir^2h)

OpenStudy (anonymous):

putting the values given

OpenStudy (anonymous):

ive done tht i got stuck solving

OpenStudy (anonymous):

see STEP 1: given that 192pi cm^3 = 1/3 ( pi * 8^2 * h)

OpenStudy (anonymous):

now 192pi cm^3 = 1/3 (64pi*h)

OpenStudy (anonymous):

192 pi cm^3 * 3 = 64pi * h

OpenStudy (anonymous):

(192 * 3)/64 cm = h 3*3 cm = h 9cm = h

OpenStudy (anonymous):

vol = 1 cm^3 ? or what ?